In a ongoing World Chess Championship, there’s a 16% chance that Tom will win the next game, and a 13% chance that Jeff will win the the next game. Whats the probability that the next will result in a draw? (In chess one of the two players wins or there is a draw- there are no more possibilities)
Answer #1

P(Tom win) = 0.16

P(Jeff win) = 0.13

Total Probability = 1

i.e.,

P(Tom win) + P(Jeff win) + P(Draw) = 1

Substituting, we get:

0.16 + 0.13 +P(Draw) = 1

So,

P(Draw) = 1- 0.29 = 0.71

So,

Answer is:

0.71